Amber would like to go on the field trip this Saturday. However, she was scheduled to
work 10 hours and she makes $8.50 an hour. Additionally, the field trip activity fee was
$30. What is her opportunity cost to go on the field trip? Can it be solved ASAP

Answer by Theo(13342)

if she goes on the field trip, it will cost her 30 dollars.
if she doesn't go on the field trip and works for 10 hours instead, she will gain 85 dollars.

by going on the field trip, it will cost her 30 dollars plus the 85 dollars she loses by not working.

her total cost for going on the field trip is 30 + 85 = 115 dollars.

the definition of opportunity cost is the return on the foregone option minus the return on the chosen option.

see this reference from https://www.investopedia.com/terms/o/opportunitycost.asp

assume she had x dollars before making the choice.
after the chosen choice she has x - 30 dollars.
after the foregone choice she has x + 85 dollars.

the opportunity cost is the foregone choice minus the chosen choice.
that would be x + 85 - (x - 30).
simplify that to get x + 85 - x + 30
combine like terms to get 115.

that, i believe, is her opportunity cost.

here's an example from the reference.

Opportunity Cost = FO − CO
where:
FO = Return on best foregone option
CO = Return on chosen option

The formula for calculating an opportunity cost is simply the difference between the expected returns of each option. Say that you have option A: to invest in the stock market hoping to generate capital gain returns. Option B, on the other hand is: to reinvest your money back into the business, expecting that newer equipment will increase production efficiency, leading to lower operational expenses and a higher profit margin.

Assume the expected return on investment in the stock market is 12 percent over the next year, and your company expects the equipment update to generate a 10 percent return over the same period. The opportunity cost of choosing the equipment over the stock market is (12% - 10%), which equals two percentage points. In other words, by investing in the business, you would forgo the opportunity to earn a higher return.

in your example, your expected return from going on the field trip is minus 30 dollars and your expected return from not going on the field trip and working instead is 85 dollars.

FO - CO becomes 85 - (-30) = 115.